d^2y/dx^2 + 2dy/dx +5y=34sinxcosx


1
Expert's answer
2021-05-19T17:49:32-0400

First we solve the related homogeneous equation 


"y''+2y'+5=0"

The characteristic equation is


"r^2+2r+5=0"

"r_1=-1-2i, r_2=-1+2i"

Hence, the general solution of the homogeneous equation is given by


"y_0(x)=C_1e^{-x}\cos 2x+C_2e^{-x}\sin 2x"

where "C_1, C_2"  are constant numbers.


Let’s go back to the nonhomogeneous equation


"y''+2y'+5y=34\sin x \cos x"

"y''+2y'+5=17\sin 2x"

Find a particular solution of the nonhomogeneous differential equation.


"y_1=(Ax+B)\cos 2x+(Cx+D)\sin 2x"

The derivatives are given by


"y_1'=A\cos 2x-2(Ax+B)\sin 2x"

"+C\sin 2x+2(Cx+D)\cos 2x"

"y_1''=-2A\sin 2x-2A\sin 2x-4(Ax+B)\cos 2x"

"+2C\cos 2x+2C\cos 2x-4(Cx+D)\sin 2x"

Substituting the function "y_1" and its derivatives in the differential equation yields:


"-2A\sin 2x-2A\sin 2x-4(Ax+B)\cos 2x"

"+2C\cos 2x+2C\cos 2x-4(Cx+D)\sin 2x"

"+2A\cos 2x-4(Ax+B)\sin 2x"

"+2C\sin 2x+4(Cx+D)\cos 2x"

"+5(Ax+B)\cos 2x+5(Cx+D)\sin 2x=17\sin 2x"



"-4A-4Cx-4D-4Ax-4B+2C=17"

"-4Ax-4B+4C+2A+4Cx+4D=0"


"-4A-4C=0"

"-4A-4B+2C-4D=17"

"-4A+4C=0"

"2A-4B+4C+4D=0"

"A=C=0"

"B=D"

"-8B=17"

Therefore we will look for a particular solution of the form


"y_1=-\dfrac{17}{8}\cos 2x-\dfrac{17}{8}\sin 2x"

Now we can write the full solution of the nonhomogeneous equation:



"y=C_1e^{-x}\cos 2x+C_2e^{-x}\sin 2x"

"-\dfrac{17}{8}\cos 2x-\dfrac{17}{8}\sin 2x"